# AUTHOR ONE et al

### v1

@@ 5 Discussion
In this paper, we revisited the question of what RR maximizes power in event-driven survival trials under PH. Our journey led us from comparing three approximations for the logrank test (Section 2), to considering other factors besides power that might influence the choice of RR (Section 3), and finally to generating insights from simulations of when unequal randomization might be of practical interest (Section 4). Here is a list of summary points:
@@
    For event-driven survival trials with HR between 0.5 and 2, Schoenfeld’s approximation remains accurate under 1:1 randomization. However, Schoenfeld’s approximation can be inaccurate under unequal randomization; it can underestimate power of the logrank test by as much as 4% absolute error, e.g., when hazard ratio and event-patient ratio are both small. Likewise, Freedman’s approximation may over or underestimate power by 3% under unequal randomization. Only Rubinstein’s approximation consistently estimates empirical power within 1% (Section 2.4).
@@
    For event-driven survival trials under PH, the RR that maximizes power of the logrank test is the RR that balances number of events across treatment arms at the end of the trial. This was shown to be true for exponential survival distributions (theoretically via Rubinstein’s approximation, empirically via Figure 3) and for more general survival distributions under PH (heuristically in Section 2.5).
@@
    3:2 and 2:1 randomization may lead to some loss of power compared to 1:1 randomization, but not to the level implied by Schoenfeld’s approximation (Figure 2). We encourage statisticians to use Rubinstein’s approximation or to perform simulations when evaluating the operating characteristics of the logrank statistic, especially for survival trials employing unequal randomization. For example, see R packages npsurvSS32 and rpact33.
@@
    Although Schoenfeld’s approximation is inaccurate under unequal randomization, we recommend that the practice of conducting event-driven survival trials to continue, with the caveat that the required number of events be carefully calculated via Rubinstein’s approximation or simulations (Section 4.2). This is because power is more sensitive to the hazard ratio compared to other design parameters such as control survival distribution and event-patient ratio. This is also consistent with current practice in survival trials that anticipate NPH at the time of trial design.
@@
    1:1 randomization may not maximize power under PH, but compared to the RR that does, the decrease in power from employing 1:1 randomization is often inconsequential (Figure 3). Likewise, compared to 3:2 and 2:1 randomization, 1:1 randomization requires a few additional events at most to maintain the same level of power (Section 4.2). All this is to say that other reasons besides power should be considered when deciding to employ unequal randomization, e.g., accrual duration, trial duration, sample size, balance in baseline characteristics, operating characteristics for estimates of secondary estimands, trial costs, individual vs. collective ethics (Section 3).
  6. 6.
    In general, 3:2 and 2:1 randomization require similar number of events as 1:1 randomization. 3:2 randomization may lead to a minor increase in trial duration if accrual rate remains the same. This increase can be mitigated by a moderate increase in accrual rate or minor increase in number of patients. In contrast, 2:1 randomization may lead to a more noticeable increase in trial duration, an increase that requires greater accrual rate or number of patients to avoid. In some cases, delay in trial readout is impossible to avoid (Section 4.2).
Point 1 has potential, important ramifications beyond sample size and power calculations under PH. Indeed, Schoenfeld’s approximation has been widely used in other settings: trial design under non-proportional hazards34, adaptive designs such as event size re-estimation35, safety monitoring of overall survival25, assurance calculation36, and cost-efficiency analysis12, to name a few. It is therefore important in these settings to understand the potential shortcomings of Schoenfeld’s approximation and to offer a more generalizable solution if necessary.
In general, sample size and power calculations are more complicated under NPH compared to under PH, because they are (even) more sensitive to design assumptions such as survival distributions and event-patient ratio. A future direction for research could be coming up with guidelines or considerations for RR under specific types of NPH, e.g., delayed treatment effect, crossing hazards, and diminishing treatment effects over time.
@@
Finally, our conclusions in Point 6 should not be interpreted by the reader as encouraging or discouraging unequal randomization. Our hope for this paper is rather to highlight considerations that trial sponsors may have missed in the past so that more accurate assessments and efficient team discussions can take place during trial design. From our collective experience, some study teams today still believe that 1:1 randomization maximizes power, and that if unequal randomization in favor of the treatment arm accelerates accrual, it will shorten the overall timeline of the trial. However—as we have argued, shown, and hopefully convinced you of through this paper—neither is necessarily true. Choosing a randomization ratio in practice requires more careful conversations, conversations in which statisticians are well-positioned to contribute to.
@@
Table S1. ‘Unequal randomization, Prolonged trial’ vs. ‘1:1’, case-by-case results.
Table S2. ‘Unequal randomization, Accelerated accrual’ vs. ‘1:1’, case-by-case results.
Table S3. ‘Unequal randomization, Increased enrollment’ vs. ‘1:1’, case-by-case results.
@@ Appendix A Ratio of expected number of events
Appendix B Piecewise exponential survival under proportional hazards
B.1 Large-sample distribution of the logrank statistic
Let $\vee$ denote the maximum and $\psi=e^{\theta}$ the hazard ratio. Furthermore, for each patient, denote his or her treatment arm by $T_{i}$ (0 = control, 1 = experimental), observed time-to-event by $X_{i}$, and censoring indicator by $\delta_{i}$ (0 = censoring, 1 = event). We assume survival in the control arm follows a piecewise exponential distribution with hazard and cumulative hazard at time $x$ equal to $\lambda_{0}(x)=\sum_{j=1}^{J}\lambda_{0j}1(t_{j-1}\leq x<t_{j})$ and $\Lambda_{0}(x)=\sum_{j=1}^{J}\lambda_{0j}[0\vee((t_{j}\wedge x)-t_{j-1})]$, respectively. Then the log-likelihood function is given by
$\ell(\psi,\lambda_{01},...,\lambda_{0J})=\sum_{i=1}^{n}\delta_{i}T_{i}\log(\psi)+\delta_{i}\log\lambda_{0}(X_{i})-\psi^{T_{i}}\Lambda_{0}(X_{i})\$
Following standard maximum likelihood theory, it can be shown that
$\frac{d\ell}{d\psi}=D_{1\cdot}/\psi-\sum_{j=1}^{J}\lambda_{0j}R_{1j}$
$\frac{d\ell}{d\lambda_{0j}}=D_{\cdot j}/\lambda_{0j}-\psi R_{1j}-R_{0j},$
where $D_{ij}$ denotes the total number of events observed in arm $i$ over time interval $j$, $D_{i\cdot}=\sum_{j=1}^{J}D_{ij}$ the total number of events observed in arm $i$, $D_{\cdot j}=\sum_{i=1}^{2}D_{ij}$ the total number of events observed over time interval $j$, and $R_{ij}=\sum_{i}1(T_{i}=i)[0\vee((t_{j}\wedge X_{i})-t_{j-1})]$ the total observed time at risk in arm $i$ over time interval $j$. Setting the first derivatives to zero, we can solve for the maximum likelihood estimates (MLEs) of $\psi$ and $\lambda_{0j}$:
$\hat{\psi}=\frac{D_{1\cdot}}{\sum_{j=1}^{J}\hat{\lambda}_{0j}R_{1j}}$
$\hat{\lambda}_{0j}=\frac{D_{0j}}{R_{0j}}.$
We have also the relationship $\hat{\psi}\hat{\lambda}_{0j}R_{1j}=D_{\cdot j}-\hat{\lambda}_{0j}R_{0j}=D_{1j}$ for all $j$.
The observed information $\mathcal{I}$ matrix can be written as
$\mathcal{I}=\begin{pmatrix}\mathcal{I}_{11}&\mathcal{I}_{12}\\ \mathcal{I}_{12}^{T}&\mathcal{I}_{22}\end{pmatrix},$
where
$\mathcal{I}_{11}=D_{1\cdot}/\psi^{2}$
$\mathcal{I}_{12}=(R_{11},...,R_{1J})$
$\mathcal{I}_{22}=\mbox{diag}(D_{\cdot 1}/\lambda_{01}^{2},...,D_{\cdot J}/\lambda_{0J}^{2}).$
We use $\mathcal{I}^{-1}$, the inverse of the observed information, to approximate the asymptotic variance of the MLEs. In particular,
$Var(\hat{\psi})=(\mathcal{I}^{-1})_{11}=(\mathcal{I}_{11}-\mathcal{I}_{12}\mathcal{I}_{22}^{-1}\mathcal{I}_{12}^{T})^{-1}=\left(\frac{D_{1\cdot}}{\psi^{2}}-\sum_{j=1}^{J}\frac{\lambda_{0j}^{2}}{D_{\cdot j}}R_{1j}^{2}\right)^{-1}.$
Plugging in the MLEs and applying the delta method, it follows that
$\widehat{Var}(\log\hat{\psi})=\left(\sum_{j=1}^{J}\left(\frac{1}{D_{1j}}+\frac{1}{D_{0j}}\right)^{-1}\right)^{-1},$
so that the wald statistic $\log\hat{\psi}/\sqrt{\widehat{Var}(\log\hat{\psi})}$ follows a normal distribution with variance 1 and mean
$\mu_{PE}=\theta\left(\sum_{j=1}^{J}\left(\frac{1}{D_{1j}}+\frac{1}{D_{0j}}\right)^{-1}\right)^{1/2}.$
We conclude that the logrank statistic follows a similar large-sample distribution since the logrank test is asymptotically efficient when there is general noninformative censoring.13 Note the difference between $\mu_{PE}$ and $\mu_{R}$ when $J=1$; $\mu_{PE}$ involves observed number of events rather than expected number of events. This is due to the fact that we used the observed information rather than the expected information to approximate the asymptotic variance of the MLEs.
B.2 Examples with a single knot ($J=2$)
Similar to Section 2.4, we first assume that dropout follows an exponential distribution such that patients have a 1% probability of dropping out every 12 months; we vary hazard ratio (HR) and event-patient ratio ($d/n$) between 0.5 and 0.8; and we calculate the target number of events $d$, sample size $n$, and accrual rate $A^{\prime}(\cdot)$ for each unique pair (HR, $d/n$). However, different from Section 2.4, we assume here that survival in the control arm from 0 to 4 months follows an exponential distribution with hazard $\lambda_{01}=\log(2)/12=0.578$, and survival in the control arm from 4 months onwards follows an exponential distribution with varying hazards $\lambda_{01}=\lambda_{00}\times\{1/2,2/3,1,3/2,2\}$.
For each scenario with a unique combination of (HR, $d/n$, $\lambda_{01}/\lambda_{00}$), we varied the randomization ratio (RR) $\phi$:1 to determine which $\phi$ minimized $\mu_{PE}$ (and hence maximized power of the logrank test). Overall, we observed at each point $\phi$ minimizing $\mu_{PE}$, that the ratio $E(D_{1\cdot})/E(D_{0\cdot})$ of expected events contributed by each arm was contained within $(0.97,1.03)$, a tight interval around 1. This supports our conclusion that, even for piecewise exponential distributions under PH, the RR that maximizes power of the logrank test is the RR that balances number of events across treatment arms at the end of the trial.
Note that having events balanced at the end of the trial does not imply that events are balanced over each time interval. Indeed, we observed at each point $\phi$ minimizing $\mu_{PE}$, that the ratio $E(D_{1j})/E(D_{0j})$ of expected events contributed by each arm over time interval $j$ varied widely. Specifically, $E(D_{11})/E(D_{01})$ varied between 0.66 and 0.94 while $E(D_{12})/E(D_{02})$ varied between 1.02 and 1.11.

### v2

@@ 5 Discussion
In this paper, we revisited the question of what RR maximizes power in event-driven survival trials under PH. Our journey led us from comparing three approximations for the LRT (Section 2), to considering other factors besides power that might influence the choice of RR (Section 3), and finally to generating insights from simulations of when 3:2 or 2:1 randomization might be of practical interest (Section 4). Here is a list of summary points:
@@
    For event-driven survival trials with HR between 0.5 and 2, Schoenfeld’s approximation remains accurate under 1:1 randomization. However, Schoenfeld’s approximation can be inaccurate under unequal randomization; it can underestimate power of the LRT by as much as 4% absolute error, e.g., when HR and EPR are both small. Likewise, Freedman’s approximation may over or underestimate power by 3%. Only Rubinstein’s approximation consistently estimates empirical power within 1% (Section 2.4).
@@
    For event-driven survival trials under PH, the RR that maximizes power of the LRT is the RR that balances number of events across treatment arms at the end of the trial. This was shown to be true for survival distributions under general PH (Sections 2.4-2.5).
@@
    1:1 randomization may not maximize power under PH, but compared to the RR that does, the decrease in power from employing 1:1 is often inconsequential (Figure 3). Likewise, compared to 3:2 and 2:1 randomization, 1:1 requires a few additional events at most to maintain the same level of power (Section 4.2). All this is to say that other reasons besides power should be considered when deciding to employ equal or unequal randomization, e.g., accrual duration, trial duration, sample size, balance in baseline characteristics, operating characteristics for estimates of secondary estimands, trial costs, individual vs. collective ethics (Section 3).
@@
    In general, 3:2 and 2:1 randomization require similar number of events as 1:1. 3:2 randomization may lead to a minor increase in trial duration if accrual rate remains the same. This increase can be mitigated by a moderate increase in accrual rate or minor increase in number of patients. In contrast, 2:1 randomization may lead to a more noticeable increase in trial duration, an increase that requires greater accrual rate or number of patients to avoid. In some cases, delay in trial readout is impossible to avoid (Section 4.2).
@@
    Although Schoenfeld’s approximation is inaccurate under unequal randomization, we recommend that the practice of conducting event-driven survival trials to continue, with the caveat that the required number of events be carefully calculated (Section 4.2). This is consistent with current practice in survival trials that anticipate NPH, even though it is known in these situations that power depends on multiple design parameters in addition to the number of events.
Point 1 has potential, important ramifications beyond sample size and power calculations under PH. Indeed, Schoenfeld’s approximation has been widely used in other settings: trial design under non-proportional hazards39, adaptive designs such as event size re-estimation40, safety monitoring of overall survival27, assurance calculation41, and cost-efficiency analysis13, to name a few. It is therefore important in these settings to understand the potential shortcomings of Schoenfeld’s approximation and to offer a more generalizable solution if necessary.
Sample size and power calculations are more complicated under NPH compared to under PH, because they are (even) more sensitive to design assumptions such as survival distributions and EPR. A future direction for research could be coming up with guidelines or considerations for RR under specific types of NPH, e.g., delayed treatment effect, crossing hazards, and diminishing treatment effects over time.
@@
Finally, Point 4 should not be interpreted by the reader as encouraging or discouraging unequal randomization. Our hope for this paper is rather to raise considerations that trial sponsors may have missed in the past so that more accurate assessments, efficient team discussions, and robust decision making can take place during trial design. For example, if there are ethical concerns in assigning patients to the control arm, then with the understanding that unequal randomization requires a similar number of events as 1:1 randomization, there might be greater incentive now to randomize patients 3:2 or 2:1. Alternatively, if there are cost and safety concerns for assigning more patients to the experimental arm than necessary, then with the understanding that unequal randomization prolongs the trial, there might be greater incentive for 1:1 randomization. Regardless of how the final decision is made, choosing a RR in practice requires careful conversations, conversations in which statisticians are well-positioned to contribute.
@@
Table S1. ‘Unequal randomization, Prolonged trial’ vs. ‘Equal randomization’, case-by-case results.
Table S2. ‘Unequal randomization, Accelerated accrual’ vs. ‘Equal randomization’, case-by-case results.
Table S3. ‘Unequal randomization, Increased enrollment’ vs. ‘Equal randomization’, case-by-case results.